# A Pairwise Comparison Relation-assisted Multi-objective Evolutionary Neural Architecture Search Method with Multi-population Mechanism

### v2

@@ 3       $acc\leftarrow$ SGD($\alpha$, $\omega_{\alpha}$)
4      $\mathcal{A}\leftarrow\mathcal{A}$cup ($\alpha$, acc)
@@ 5      $i\leftarrow i+1$
7while $t<T$ do
       $predictor\leftarrow$ Construct surrogate model from $\mathcal{A}$// Online update the $predictor$.
@@
9      $Solutions\leftarrow$ MP-MOEA($\mathcal{A}$, $predictor$, $f_{c}$)
10      $R_{t}\leftarrow$ Select $K$ elite architectures from $Solutions$
@@ 13            $acc\leftarrow$ SGD($\alpha$, $\omega_{\alpha}$)
14            $\mathcal{A}\leftarrow\mathcal{A}$cup ($\alpha$, acc)
@@ III-A Overall Framework
Algorithm 1 shows the pseudo code of the proposed algorithm. Firstly, in order to construct an efficient surrogate model, some evaluated architectures are needed to construct the training dataset. So before the search process starts, an empty archive $\mathcal{A}$ for storing training samples is initialized (line 3). $N$ individuals are randomly sampled from the search space and then are decoded for training. As mentioned above, we employ the weight sharing during the training. The weights inherited from the trained supernet are used as a warm up for the stochastic gradient descent (SGD) algorithm to improve the search efficiency (lines 4-10). Afterwards, there is a main search loop of in the algorithm (line 11). The surrogate model is constructed using the archive $\mathcal{A}$ (line 12). A detailed description of the surrogate model is given later in subsection III-C. Then we define a set to store individuals to be evaluated. Next, we use the proposed multi-objective evolutionary algorithm with multi-population mechanism (MP-MOEA for short) combine with our surrogate model to optimize both accuracy (Acc) and MAdds, i.e., maximize classification accuracy and minimize model complexity (line 13). This process is described in detail later in subsection III-D. To improve the sample efficiency of our search, we train the surrogate model using an online learning approach. The retained individuals of each generation are decoded and trained, and these retained offspring are used as new training samples to update the surrogate model (lines 15-19). The above steps are repeated until the conditions for the end of the procedure are satisfied. Finally, after the multi-objective evolutionary search, the optimal architectures are selected from the pool of architectures based on the non-dominated sorting (line 22).
@@ III-B Search Space and Encoding
The search process of NAS starts with constructing a search space that can contain most CNNs for image tasks. In this paper, the backbone structure is based on MobileNetV3 [52]. The structure consists of three stages. The start stage extracts features, and the final stage outputs categories. These two parts do not need to be searched. The middle stage needs to be searched and consists of five sequentially connected MBConvBlocks [51] that progressively decrease the feature map size and increase the number of channels. Each block is composed of a series of layers, and every layer adopts the inverted bottleneck structure [53]: first, a $1\times 1$ convolution, which is used to convert the input channel to the expansion channel; second, a depth-wise convolution, which is the expansion channel and contains the parameter stride; at last, a $1\times 1$ convolution, which is used to convert from the expansion channel to the output channel. The number of layers of each convolution block (depth) is selected from $\left\{2,3,4\right\}$; for each layer, we search the expansion rate of the first $1\times 1$ convolution and the kernel size of the depth-wise separable convolution. The expansion rate is selected from $\left\{3,4,6\right\}$, and the kernel size is selected from $\left\{3,5,7\right\}$. Moreover, we also allow the CNNs to obtain multiple input image size (resolution), ranging from 192 to 256 with a stride of 4. Therefore, there are approximately $10^{20}$ different architectures in the search space. For the encoding strategy, we use a fixed 46-bit integer string. If the encoding length of the architecture with fewer layers is less than 46 bits, the fixed length is achieved by padding with zeros. Fig. 2 shows the encoding strategy, and Fig. 3 shows the search space of this paper. As shown in these figures, the architecture is composed of five blocks. The encoding of this architecture is composed of image resolution and other parts representing five blocks. For each block, its encoding consists of the number of layers, the expansion rate, and the kernel size of each layer. In addition, the values in the encoding are all the index of considered options. As depicted in Fig. 2, the first bit “0” denotes the index of the first candidate resolution which means the resolution is 192. The remaining 45 bits represent five blocks, each of which is represented by 9 bits. e.g., the Block 1 can be encoded as “211200012”. The first bit “2” denotes the index of candidate layers which means the Block 1 has four layers. Then, the next four bits “1120” denote the kernel size of each layer. Specifically, the kernel size in each layer in order is $\left\{5,5,7,3\right\}$. Similarly, the rest four bits denote that the expansion rate in each layer in order is $\left\{3,3,4,6\right\}$. In Block 5, the gray bit “0” represents no layer and the encoding is padded with zeros, which is not from considered options.
@@ III-C Surrogate Model
Unlike most of the existing surrogate-assisted methods used to predict the classification accuracy of architectures, we use the surrogate model to obtain the performance ranking of candidate architectures. In the evolutionary algorithm, the selection operation needs to be carried out according to the fitness values of individuals, and individuals with high fitness will be selected for the subsequent evolution. Therefore, it is more important to obtain the ranking of the candidates than to predict the true accuracy. We propose a surrogate model based on pairwise comparison relations, which is constructed on the basis of a classification model. Its inputs are vectors of the encoding combinations of pairwise architectures, and its outputs are labels indicating which one is the better solution. In this paper, we try to study the comparison relation between two individuals and transform the comparison relation learning problem into a binary classification problem. Binary classification models can be trained better with fewer samples. For example, given $n$ labeled architectures, we generate training samples by comparing all unique pairs of architectures, and $\frac{n(n-1)}{2}$ training samples can be obtained. Given a small number of samples, the surrogate model with better performance can be obtained through the extended training samples.
@@
In this paper, SVM is used as the surrogate model. In order to train the surrogate model, we should firstly construct the training dataset. The construction method is the same as that proposed in EffPNet [47]. The specific process is as follows: firstly, the data used to train needs to be built from the archive $\mathcal{A}$. The $n$ individuals in the archive $\mathcal{A}$ are matched in pairs sequentially. In a pair of individuals, if the fitness value of the first individual is better than the other, the class label will be 1. Otherwise, the class label will be 0. Then $n(n-1)/2$ training samples can be obtained to feed into the SVM for training. The above data processing transforms the direct prediction of architecture performance into a binary classification task comparing the performance of a pair of architectures. We take a pair of encodings of two architectures as the input of the surrogate model, and the output is label “1” or “0” as described above. To make it more intuitive, an example is provided in Fig. 4 to present the concatenation of the encodings of two architectures, as well as the output. Label “1” indicates that architecture $Arc\_1$ has better performance, while label “0” indicates that architecture $Arc\_2$ has better performance.
@@
Through the above pairwise relations prediction, we can obtain the performance ranking of all architectures. The details are as follows: assuming that there are $n$ architectures, we set a score for each architecture corresponding to its performance. Firstly, we match the first architecture with the other $n-1$ architectures and use the trained surrogate model to predict the relationship of these pairs of architectures. For each pair of architectures, if the first architecture outperforms the second one, the score of the first architecture is added by 1, otherwise the score of the second architecture is added by 1. Then, the second architecture is paired one by one with the remaining architectures after removing the first architecture, and the above process is repeated until the last architecture is left. Finally, each architecture is compared $n-1$ times and the score ranges from 0 to $n-1$. The number of comparisons in total is $n(n-1)/2$. The larger the value of the score, the more times the corresponding architecture wins in the overall comparisons, which implies that the architecture has a superior performance. The results from pairwise relations prediction will be used in the non-dominated sorting for the environment selection during the subsequent evolutionary process.
@@ III-D Multi-objective Evolutionary Algorithm Based on Multi-population Mechanism
In multi-objective evolutionary algorithms, diversity needs to be considered as well as convergence. During the practical evolution process, it is easy to fall into the local optimum without exploring the region of the optimal solution. To overcome this problem, we propose a multi-objective evolutionary algorithm based on the multi-population mechanism, see Algorithm 2 for details. The multi-population mechanism is mainly used in the selection operation to select parents for the next generation during the evolution process. Firstly, two sub-populations are constructed from the initial population $P_{t}$ (line 1). The main population is $E_{t}$, and the vice population is $F_{t}$. Specifically, the initial population $P_{t}$ is non-dominated sorted according to the predicted architectural rankings obtained from the above surrogate model and the calculated MAdds. After that, the non-dominated solutions of the first level are taken as the main population $E_{t}$ (lines 2-5). Then, the individuals in the population $E_{t}$ are removed from the population $P_{t}$, and the crowding distances of the remaining solutions are calculated. $K$ individuals with high crowding ranking are selected to form the vice population $F_{t}$ (line 6). We set a threshold based on the number of generations, and when this threshold is less than $\theta$, Both parent $p_{1}$ and $p_{2}$ come from the main population $E_{t}$ (line 13). If this threshold is greater than $\theta$, parent $p_{1}$ comes from the primary population and parent $p_{2}$ comes from the secondary population (line 15). A number of offspring architectures are generated until the number of offspring is greater than or equal to $m$ in the current generation. $m$ is a number used to limit the number of offspring. The threshold is a value related to the evolution generation, which is generated by the following equation:
$Threshold=$$\left\{\begin{array}[]{lc}random(\delta,0.7)&g<\frac{1}{4}G\\ random(0,\delta)&\frac{1}{4}G\leq g\leq\frac{3}{4}G\\ random(\delta,1)&g>\frac{3}{4}G\\ \end{array}\right.$ (4)
$\delta=$$\left\{\begin{array}[]{lc}random(0,0.7)&g<\frac{1}{4}G\\ random(0,1)&g\geq\frac{1}{4}G\\ \end{array}\right.$ (7)
@@
Input: Initial population $P_{t}$, surrogate model $predictor$, complexity objective $f_{c}$, number of generations $G$, vice population size $K$.
@@
Predict the accuracy ranking of architectures in $P_{t}$ by $predictor$ // $predictor$ see Algo.1.
@@ 27       $g\leftarrow g+1$
TABLE I: Comparison with state-of-the-art image classifiers on the CIFAR-10 dataset. The multi-objectives used for architecture optimization are performance and model complexity. *: MAdds computed using the genotypes provided by the authors. $\dagger$ denotes results from the reimplementation. SMEM-NAS(w/o): Architectures found by SMEM-NAS without the multi-population mechanism. Search cost: GPU days of the searching process.
@@
EfficientNet-B0 [13] 98.1 387 4.0 - - 2019 manual
iDARTS [54] 97.75 - 3.6 0.2 Tesla V100 2023 GD
CDARTS [55] 97.52 - 3.9 0.3 Tesla V100 2023 GD
SWD-NAS [56] 97.49 519 3.17 0.13 RTX 3090 2024 GD
PA-DARTS [57] 97.59 578 3.75 0.36 RTX 3090 2024 GD
NoisyDARTS [58] 97.63 534 3.3 0.4 Tesla V100 2021 GD
GENAS [59] 97.55 504 3.53 0.26 RTX 3090 2024 GD
RelativeNAS [60] 97.66 - 3.93 0.4 GTX 1080Ti 2023 GD
DARTS-PT-CORE [61] 97.5 - 3.6 0.8 Tesla V100 2024 GD
NPENAS-NP [62] 97.46 - 3.5 1.8 RTX 2080Ti 2023 EA
EEEA-Net-C [63] 97.54 - 3.6 0.52 RTX 2080Ti 2021 EA
CGP-NAS [64] 96.3 636 4.04 11.5 RTX 1080Ti 2023 EA
SMCSO [65] 97.12 - 3.46 1.32 RTX 2080Ti 2025 EA
SPNAS [66] 98.2 - 6.33 1.4 RTX 3090 2025 EA
FairNAS-A [49] 98.2 391 - 12 Tesla V100 2021 EA
CARS [37] 97.38 728 3.6 0.4 Tesla V100 2020 EA
ESENet [67] 96.44 540 4.53 9 RTX 2080Ti 2023 EA
NSGANetV2 [25] 98.4 468 - - - 2020 EA
M2M-Net [68] 97.56 - 3.79 6 RTX 3080 2024 EA
NASNet-A [7] 97.35 $608^{\ast}$ 3.3 1800 Tesla P100 2018 RL
BNAS [69] 97.03 - 4.7 0.19 - 2022 RL
DBNAS [70] 97.58 - 2.4 0.9 RTX 3090 2025 RL
@@
TABLE II: Comparison with state-of-the-art image classifiers on the CIFAR-100 dataset. $\dagger$ denotes results from the reimplementation. SMEMNAS-transfer: the architecture, SMEMNAS-L, which searched on CIFAR-10.
Architecture
Top-1
MAdds
Params
Search
Cost
Device Year
Search
Method
EfficientNet-B0 [13] 88.1 400 4.0 - - 2019 manual
MobileNetV2 [53] 80.8 300 3.4 - - 2018 manual
PC-DARTS [71] 83.1 - 3.6 0.3 GTX 1080Ti 2019 GD
SWD-NAS [56] 83.87 598 3.56 0.13 RTX 3090 2024 GD
DARTS-PT-CORE [61] 83.27 - 3.6 0.8 Tesla V100 2024 GD
PA-DARTS [57] 83.62 578 3.75 0.36 RTX 3090 2024 GD
RelativeNAS [60] 84.14 - 3.98 0.4 GTX 1080Ti 2023 GD
GENAS [59] 83.14 504 3.53 0.26 RTX 3090 2024 GD
SMCSO [65] 80.66 - 3.72 2 RTX 2080Ti 2025 EA
FairNAS-A [49] 87.3 391 - 12 Tesla V100 2021 EA
CGP-NAS [64] 79.37 827 5.9 11.28 RTX 1080Ti 2023 EA
ESENet [67] 76.35 540 4.53 9 RTX 2080Ti 2023 EA
M2M-Net [68] 84.77 - 3.83 6 RTX 3080 2024 EA
SPNAS [66] 87.26 - 6.7 1.6 RTX 3090 2025 EA
NASNet-A Large [7] 86.7 12031 - 1800 Tesla P100 2018 RL
DBNAS [70] 85.18 - 2.5 0.9 RTX 3090 2025 RL
SMEMNAS-S(w/o) 86.05 194 3.5 1.25 RTX 3090 - EA
SMEMNAS-M(w/o) 86.4 242 4.1 1.25 RTX 3090 - EA
SMEMNAS-L(w/o) 86.65 348 5.8 1.25 RTX 3090 - EA
SMEMNAS-transfer 87.18 305 5.0 1.25 RTX 3090 - EA
SMEMNAS-S(ours) 86.71 197 3.9 1.25 RTX 3090 - EA
SMEMNAS-M(ours) 87.25 250 4.1 1.25 RTX 3090 - EA
SMEMNAS-L(ours) 87.95 755 6.4 1.25 RTX 3090 - EA
IV Experiments
In this section, we conducted a series of experiments to validate the effectiveness of our proposed method SMEM-NAS. Specifically, we firstly show our experimental configurations in Section IV-A. Next, Section IV-B presents and analyses our experimental results on benchmark datasets. Then we evaluate the effectiveness of the multi-population mechanism in Section IV-C. The comparison results with other algorithms in terms of different perspectives are discussed, including accuracy, complexity, search cost, device and year of publication. Finally, Section IV-D shows some ablation experiments for surrogate models and the initial population.
IV-A Experimental Configurations
For each dataset, we started with 100 randomly sampled architectures from the search space as the initial population. The evolutionary search process was executed for a total of 25 iterations. Except for the initial population, for each subsequent iteration, we retained and evaluated eight architectures from generated offspring. They were used for the training of the surrogate model and for the next iteration. So, in the whole search process, we evaluated 300 architectures in total, which were also used to train the surrogate model. Since evaluating architecture performance is time-time-consuming, we adopted the weight-sharing technique to accelerate the training process to improve the search efficiency. The OnceForAll [51] supernet is trained with the progressive shrinking algorithm, in which the largest network is fine-tuned to support sub-networks and they are added to the sampling space by sharing weights. During the search process, the architecture weights were inherited from the trained OnceForAll as initial weights for training. Because OnceForAll is based on the ImageNet dataset, we fine-tuned the weights inherited from the supernet for five epochs when searching on CIFAR-10 and CIFAR-100. All our experiments were performed on a single Nvidia RTX 3090 GPU card. Finally, we selected three promising architectures from the obtained non-dominated solutions and then trained them for 200 epochs by the standard SGD optimizer with momentum, where the initial learning rate, the momentum rate and the batch size were set to 0.01, 0.9 and 128. More detailed information can be found in our publicly available code https://github.com/ccz-enas/SMEM-NAS.

### v3

@@ 3 $acc\leftarrow$ SGD($\alpha$, $\omega_{\alpha}$)
4 $\mathcal{A}\leftarrow\mathcal{A}$ $\cup$ ($\alpha$, acc)
@@ 5 $i\leftarrow i+1$
7 while $t<T$ do
 $predictor\leftarrow$ Construct surrogate model from $\mathcal{A}$// Online update the predictorpredictor.
@@
9 $Solutions\leftarrow$ MP-MOEA($\mathcal{A}$, predictorpredictor, $f_{c}$)
10 $R_{t}\leftarrow$ Select $K$ elite architectures from SolutionsSolutions
@@ 13    $acc\leftarrow$ SGD($\alpha$, $\omega_{\alpha}$)
14    $\mathcal{A}\leftarrow\mathcal{A}$ $\cup$ ($\alpha$, acc)
@@ III-A Overall Framework
Algorithm 1 shows the pseudo code of the proposed algorithm. Firstly, in order to construct an efficient surrogate model, some evaluated architectures are needed to construct the training dataset. So before the search process starts, an empty archive $\mathcal{A}$ for storing training samples is initialized (line 1). $N$ individuals are randomly sampled from the search space and then are decoded for training. As mentioned above, we employ the weight sharing during the training. The weights inherited from the trained supernet are used as a warm-up for the stochastic gradient descent (SGD) algorithm to improve the search efficiency (lines 1-1). Afterwards, there is a main search loop of in the algorithm (line 1). The surrogate model is constructed using the archive $\mathcal{A}$ (line 1). A detailed description of the surrogate model is given later in subsection III-C. Then we define a set to store individuals to be evaluated. Next, we use the proposed multi-objective evolutionary algorithm with multi-population mechanism (MP-MOEA for short) combine with our surrogate model to optimize both accuracy (Acc) and MAdds, i.e., maximize classification accuracy and minimize model complexity (line 1). This process is described in detail later in subsection III-D. To improve the sample efficiency of our search, we train the surrogate model using an online learning approach. The retained individuals of each generation are decoded and trained, and these retained offspring are used as new training samples to update the surrogate model (lines 1-1). The above steps are repeated until the conditions for the end of the procedure are satisfied. Finally, after the multi-objective evolutionary search, the optimal architectures are selected from the pool of architectures based on the non-dominated sorting (line 1).
@@ III-B Search Space and Encoding
The search process of NAS starts with constructing a search space that can contain most CNNs for image tasks. In this paper, the backbone structure is based on MobileNetV3 [howard_searching_2019]. The structure consists of three stages. The start stage extracts features, and the final stage outputs categories. These two parts do not need to be searched. The middle stage needs to be searched and consists of five sequentially connected MBConvBlocks [cai_once_2020] that progressively decrease the feature map size and increase the number of channels. Each block is composed of a series of layers, and every layer adopts the inverted bottleneck structure [sandler_mobilenetv2_2018]: first, a $1\times 1$ convolution, which is used to convert the input channel to the expansion channel; second, a depth-wise convolution, which is the expansion channel and contains the parameter stride; at last, a $1\times 1$ convolution, which is used to convert from the expansion channel to the output channel. The number of layers of each convolution block (depth) is selected from $\left\{2,3,4\right\}$; for each layer, we search the expansion rate of the first $1\times 1$ convolution and the kernel size of the depth-wise separable convolution. The expansion rate is selected from $\left\{3,4,6\right\}$, and the kernel size is selected from $\left\{3,5,7\right\}$. Moreover, we also allow the CNNs to obtain multiple input image size (resolution), ranging from 192 to 256 with a stride of 4. Therefore, there are approximately $10^{20}$ different architectures in the search space. For the encoding strategy, we use a fixed 46-bit integer string. If the encoding length of the architecture with fewer layers is less than 46 bits, the fixed length is achieved by padding with zeros. Fig. 2 shows the encoding strategy, and Fig. 3 shows the search space of this paper. As shown in these figures, the architecture is composed of five blocks. The encoding of this architecture is composed of image resolution and other parts representing five blocks. For each block, its encoding consists of the number of layers, the expansion rate, and the kernel size of each layer. In addition, the values in the encoding are all the index of considered options. As depicted in Fig. 2, the first bit “0” denotes the index of the first candidate resolution which means the resolution is 192. The remaining 45 bits represent five blocks, each of which is represented by 9 bits. e.g., the Block 1 can be encoded as “211200012”. The first bit “2” denotes the index of candidate layers which means the Block 1 has four layers. Then, the next four bits “1120” denote the kernel size of each layer. Specifically, the kernel size in each layer in order is $\left\{5,5,7,3\right\}$. Similarly, the rest four bits denote that the expansion rate in each layer in order is $\left\{3,3,4,6\right\}$. In Block 5, the gray bit “0” represents no layer and the encoding is padded with zeros, which is not from considered options.
@@ III-C Surrogate Model
Unlike most of the existing surrogate-assisted methods used to predict the classification accuracy of architectures, we use the surrogate model to obtain the performance ranking of candidate architectures. In the evolutionary algorithm, the selection operation needs to be carried out according to the fitness values of individuals, and individuals with high fitness will be selected for the subsequent evolution. Therefore, it is more important to obtain the ranking of the candidates than to predict the true accuracy. We propose a surrogate model based on pairwise comparison relations, which is constructed on the basis of a classification model. Its inputs are vectors of the encoding combinations of pairwise architectures, and its outputs are labels indicating which one is the better solution. In this paper, we try to study the comparison relation between two individuals and transform the comparison relation learning problem into a binary classification problem. Binary classification models can be trained better with fewer samples. For example, given $n$ labeled architectures, we generate training samples by comparing all unique pairs of architectures, and n(n1)/2n(n-1)/2 training samples can be obtained. Given a small number of samples, the surrogate model with better performance can be obtained through the extended training samples.
@@
In this paper, SVM is used as the surrogate model. In order to train the surrogate model, we should firstly construct the training dataset. The construction method is the same as that proposed in EffPNet [wang_surrogate-assisted_2022]. The specific process is as follows: firstly, the data used to train needs to be built from the archive $\mathcal{A}$. The $n$ individuals in the archive $\mathcal{A}$ are matched in pairs sequentially. In a pair of individuals, if the fitness value of the first individual is better than the other, the class label will be 1. Otherwise, the class label will be 0. Then n(n1)/2n(n-1)/2 training samples can be obtained to feed into the SVM for training. The above data processing transforms the direct prediction of architecture performance into a binary classification task comparing the performance of a pair of architectures. We take a pair of encodings of two architectures as the input of the surrogate model, and the output is label “1” or “0” as described above. To make it more intuitive, an example is provided in Fig. 4 to present the concatenation of the encodings of two architectures, as well as the output. Label “1” indicates that architecture Arc_1Arc\_1 has better performance, while label “0” indicates that architecture Arc_2Arc\_2 has better performance.
@@
Through the above pairwise relations prediction, we can obtain the performance ranking of all architectures. The details are as follows: assuming that there are $n$ architectures, we set a score for each architecture corresponding to its performance. Firstly, we match the first architecture with the other n1n-1 architectures and use the trained surrogate model to predict the relationship of these pairs of architectures. For each pair of architectures, if the first architecture outperforms the second one, the score of the first architecture is added by 1, otherwise the score of the second architecture is added by 1. Then, the second architecture is paired one by one with the remaining architectures after removing the first architecture, and the above process is repeated until the last architecture is left. Finally, each architecture is compared n1n-1 times and the score ranges from 0 to n1n-1. The number of comparisons in total is n(n1)/2n(n-1)/2. The larger the value of the score, the more times the corresponding architecture wins in the overall comparisons, which implies that the architecture has a superior performance. The results from pairwise relations prediction will be used in the non-dominated sorting for the environment selection during the subsequent evolutionary process.
@@ III-D Multi-objective Evolutionary Algorithm Based on Multi-population Mechanism
In multi-objective evolutionary algorithms, diversity needs to be considered as well as convergence. During the practical evolution process, it is easy to fall into the local optimum without exploring the region of the optimal solution. To overcome this problem, we propose a multi-objective evolutionary algorithm based on the multi-population mechanism, see Algorithm 2 for details. The multi-population mechanism is mainly used in the selection operation to select parents for the next generation during the evolution process. Firstly, two sub-populations are constructed from the initial population $P_{t}$ (line 2). The main population is $E_{t}$, and the vice population is $F_{t}$. Specifically, the initial population $P_{t}$ is non-dominated sorted according to the predicted architectural rankings obtained from the above surrogate model and the calculated MAdds. After that, the non-dominated solutions of the first level are taken as the main population $E_{t}$ (lines 2-2). Then, the individuals in the population $E_{t}$ are removed from the population $P_{t}$, and the crowding distances of the remaining solutions are calculated. $K$ individuals with high crowding ranking are selected to form the vice population $F_{t}$ (line 2). We set a threshold based on the number of generations, and when this threshold is less than $\theta$, Both parent $p_{1}$ and $p_{2}$ come from the main population $E_{t}$ (line 2). If this threshold is greater than $\theta$, parent $p_{1}$ comes from the primary population and parent $p_{2}$ comes from the secondary population (line 2). A number of offspring architectures are generated until the number of offspring is greater than or equal to $m$ in the current generation. $m$ is a number used to limit the number of offspring. The threshold is a value related to the evolution generation, which is generated by the following equation:
[…]
@@
Input: Initial population $P_{t}$, surrogate model predictorpredictor, complexity objective $f_{c}$, number of generations $G$, vice population size $K$.
@@
Predict the accuracy ranking of architectures in $P_{t}$ by predictorpredictor // predictorpredictor see Algo.1.
@@ 27 $g\leftarrow g+1$
IV Experiments
In this section, we conducted a series of experiments to validate the effectiveness of our proposed method SMEMNAS. Specifically, we firstly show our experimental configurations in Section IV-A. Next, Section IV-B presents and analyses our experimental results on benchmark datasets. Then we evaluate the effectiveness of the multi-population mechanism in Section IV-C. The comparison results with other algorithms in terms of different perspectives are discussed, including accuracy, complexity, search cost, device and year of publication. Finally, Section IV-D shows some ablation experiments for surrogate models and the initial population.
IV-A Experimental Configurations
For each dataset, we started with 100 randomly sampled architectures from the search space as the initial population. The evolutionary search process was executed for a total of 25 iterations. Except for the initial population, for each subsequent iteration, we retained and evaluated eight architectures from generated offspring. They were used for the training of the surrogate model and for the next iteration. So, in the whole search process, we evaluated 300 architectures in total, which were also used to train the surrogate model. Since evaluating architecture performance is time-consuming, we adopted the weight-sharing technique to accelerate the training process and improve the search efficiency. The Once-For-All [cai_once_2020] supernet is trained with the progressive shrinking algorithm, in which the largest network is fine-tuned to support sub-networks and they are added to the sampling space by sharing weights. During the search process, the architecture weights were inherited from the trained Once-For-All as initial weights for training. Because Once-For-All is based on the ImageNet dataset, we fine-tuned the weights inherited from the supernet for five epochs when searching on CIFAR-10 and CIFAR-100. All our experiments were performed on a single Nvidia RTX 3090 GPU card. Finally, we selected three promising architectures from the obtained non-dominated solutions and then trained them for 200 epochs by the standard SGD optimizer with momentum, where the initial learning rate, the momentum rate and the batch size were set to 0.01, 0.9 and 128. More detailed information can be found in our publicly available code at https://github.com/pcjiang1998/SMEM-NAS.
TABLE I: Comparison with state-of-the-art image classifiers on the CIFAR-10 dataset. The multi-objectives used for architecture optimization are performance and model complexity. SMEMNAS(w/o): Architectures found by SMEMNAS without the multi-population mechanism. Search cost: GPU days of the searching process.
@@ IV-A Experimental Configurations
EfficientNet-B0 [tan_efficientnet_2019] 98.1 387 4.0 - - 2019 manual
iDARTS [wang_idarts_2023] 97.75 - 3.6 0.2 Tesla V100 2023 GD
CDARTS [yu_cyclic_2023] 97.52 - 3.9 0.3 Tesla V100 2023 GD
SWD-NAS [xue_self-adaptive_2024] 97.49 519 3.17 0.13 RTX 3090 2024 GD
PA-DARTS [xue_improved_2024] 97.59 578 3.75 0.36 RTX 3090 2024 GD
GENAS [xue_gradient-guided_2024] 97.55 504 3.53 0.26 RTX 3090 2024 GD
RelativeNAS [tan_relativenas_2023] 97.66 - 3.93 0.4 GTX 1080Ti 2023 GD
NPENAS-NP [wei_npenas_2023] 97.46 - 3.5 1.8 RTX 2080Ti 2023 EA
CGP-NAS [garcia-garcia_continuous_2023] 96.3 636 4.04 11.5 RTX 1080Ti 2023 EA
SMCSO [xue_surrogate_2025] 97.12 - 3.46 1.32 RTX 2080Ti 2025 EA
SPNAS [jiang_score_2025] 98.2 - 6.33 1.4 RTX 3090 2025 EA
FairNAS-A [chu_fairnas_2021] 98.2 391 - 12 Tesla V100 2021 EA
CARS [yang_cars_2020] 97.38 728 3.6 0.4 Tesla V100 2020 EA
ESENet [qiu_efficient_2023] 96.44 540 4.53 9 RTX 2080Ti 2023 EA
NSGANetV2 [lu_nsganetv2_2020] 98.4 468 - - - 2020 EA
M2M-Net [tan_m2m-net_2024] 97.56 - 3.79 6 RTX 3080 2024 EA
NASNet-A [zoph_learning_2018] 97.35 608 3.3 1800 Tesla P100 2018 RL
BNAS [ding_bnas_2022] 97.03 - 4.7 0.19 - 2022 RL
DBNAS [yang_deeply_2025] 97.58 - 2.4 0.9 RTX 3090 2025 RL